Bacterial motion. A bacterium using its flagellum as propulsion can move through liquids at a rate of 0.003 m/s. For a 50-?m long bacterium, that is the equivalent of 60 cell lengths per second. Bacteria of that size have a mass of approximately 1
Answer #1

a = v^2-u^2/2s = -0.003*0.003/(2*0.01*10^-9) = 450000 m/s^2

F = ma = 4.5*10^-10 N